OpenStudy (anonymous):

find the value of the discriminant of the quadratic equation and comment on the nature of the roots. Solve the equation. 4x2 – 17x = 15 please show work.

OpenStudy (saifoo.khan):

b^2 - 4ac a = 4 b = -17 c = -15 (-17)^2 - 4(4)(-15) 529

OpenStudy (saifoo.khan):

as the result we got is Positive so it means that we will have 2 real roots.

OpenStudy (anonymous):

and since by some miracle 529 is a perfect square (it is the square of 23) your two solutions will be rationals as well

OpenStudy (anonymous):

\[\sqrt{b ^{2}-4ac}\] So we have \[4x^{2}-17x-15=0\] \[\sqrt{(-17) ^{2}-4(4)(-15)}=529\] 529 = 23^2, So there will be two rational roots.